A 600 g ball A is moving with a velocity of magnitude 6 m/s when it hit by a 1 kg ball B that has a velocity of magnitude 4 m/s. Knowing that the coefficient of restitution is 0.8 and assuming no friction, determine the velocity of each ball after impact.
1
Expert's answer
2020-07-22T06:03:36-0400

"(v_A)_n=6\cdot \cos40°=4.596m/s"


"(v_A)_t=-6\cdot \sin40°=-3.857m/s"


"(v_B)_n=-4m\/s"


"(v_B)_t=0"


t-direction:


"m_A(v_A)_t+m_B(v_B)_t=m_A(v'_A)_t+m_B(v'_B)_t"


"0.6\cdot(-3.857)+0=0.6\cdot(v'_A)_t+1\cdot (v'_B)_t"


"-2.314=0.6\cdot(v'_A)_t+(v'_B)_t"


For ball A


"m_A\cdot(v_A)_t=m_A\cdot(v'_A)_t\to -3.857=(v'_A)_t\to (v'_A)_t=-3.857m/s"


So, we have


"-2.314=0.6\cdot(-3.857)+(v'_B)_t\to (v'_B)_t=0"


n-direction


"((v_A)_n-(v_B)_n)e=(v'_B)_n-(v'_A)_n"

"(4.596-(-4))\cdot0.8=(v'_B)_n-(v'_A)_n"


"(v'_B)_n-(v'_A)_n=6.877\to (v'_B)_n=6.877+(v'_A)_n"


"m_A(v_A)_n+m_B(v_B)_n=m_A(v'_A)_n+m_B(v'_B)_n"


"0.6\cdot 4.596+1\cdot(-4)=1\cdot(v'_B)_n+0.6\cdot(v'_A)_n"


"(v'_B)_n+0.6\cdot(v'_A)_n=-1.2424\to 6.877+(v'_A)_n+0.6\cdot(v'_A)_n=-1.2424"


"(v'_A)_n=-5.075m\/s"


"(v'_B)_n=6.877+(v'_A)_n=6.877-5.075=1.802m\/s"


For ball A after impact


"\tan\beta=\frac{(v_A)_t}{(v_A)_n}=\frac{3.857}{5.075}\to \beta=37.2°\to \beta+40°=77.2°" Answer


"v'_A=\sqrt{3.857^2+5.075^2}=6.37m/s" Answer


For ball B after impact


"v'_B=1.802m\/s", "\gamma=40°" Answer